# Language Models That Walk the Talk: A Framework for Formal Fairness Certificates

### v1

@@ C.3 Ablation studies
C.3.1 Fairness scores without pre-training the Keras embedding layer
@@
The results for the 50 nearest neighbors and their corresponding distances for the words female and male are presented in Table 11, respectively. As shown, in the absence of pre-training, the nearest neighbors seldom reflect strongly gendered semantics. This underscores the importance of embedding layer pre-training in capturing meaningful semantic relationships, thereby justifying the design of our experimental framework. For comparison, the distance results with pre-training are provided in Section C.3.4.
For the sake of completeness, we conducted the fairness score assessment despite the absence of gendered words among the 50 nearest neighbors for both gender indicators, we simplified the fairness evaluation by setting the threshold based solely on the $\ell_{\infty}$ distance between female and male (which is the gender indicator in the Adult Census dataset), as the reference circle (which is 0.7581), rather than using the maximum distance among the 50 neighboring words of the two gender indicators. Please note that the entire experimental setup, including the language model training parameters, remains identical to the experiment that includes the pre-training of the embedding model (Section 4.2). The results of the formal verification for the model without any pre-training of the Keras embedding layer are presented in Figure 6. As shown, not only do all sentences fail verification, but their corresponding $\ell_{\infty}$ radii also lie significantly within the reference circle.
Table 11: Top 50 most $\ell_{\infty}$similar words to female and male using the Keras embedding layer without pre-training.
Model Salary Prediction & Gender Bias Mitigation (8 blocks)
@@
Word Dist. Word Dist. Word Dist. Word Dist.
@@
7th $1.488\text{e}{-01}$ 6th $1.504\text{e}{-01}$#ec $1.465\text{e}{-01}$ specialty $1.474\text{e}{-01}$
5th $1.560\text{e}{-01}$ 8 $1.672\text{e}{-01}$ 85 $1.477\text{e}{-01}$#em $1.519\text{e}{-01}$
1st $1.753\text{e}{-01}$ 5 $1.794\text{e}{-01}$ 13 $1.534\text{e}{-01}$#v $1.568\text{e}{-01}$
@@
dominican $2.369\text{e}{-01}$#vi $2.384\text{e}{-01}$ england $2.080\text{e}{-01}$ lee $2.130\text{e}{-01}$
@@
C.3.2 Relationship between embedding space and max perturbed noise radius
This section investigates the relationship between the margin parameter in (2), the resulting embedding bounds, and the maximum verifiable perturbation radius. As the margin parameter decreases, the embedding bounds contract accordingly, leading to a reduction in the verifiable radius. This contraction can impede the successful verification of semantically similar words, as their embedding distances may fall outside the increasingly tightened bounds. An illustrative graph Figure 7 highlights the trade-off between margin size and verifiability within semantic clusters [^12] The range in the graph is for demonstrative purposes only; for more detailed metrics corresponding to specific margin values, please refer to Section C.2..
@@ C.3.2 Relationship between embedding space and max perturbed noise radius
C.3.3 Transformer-neighbor configurations
In this section, we examine how fairness scores are influenced by varying combinations of transformer block counts and the number of nearest neighbors. The experimental configuration for this analysis is illustrated in Figure 8. As previously discussed in Section 4.2, we presented fairness scores for selected models with different transformer depths. Here, our objective is to provide a more comprehensive view, highlighting how fairness metrics evolve progressively as we adjust the number of transformer blocks and nearest neighbors. This analysis aims to uncover systematic trends and sensitivities within the model architecture that impact fairness.
@@ C.3.3 Transformer-neighbor configurations
We illustrate the progressive changes in fairness scores for the 6-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 17, 17 and 17131313The selected numbers of neighbors are intended to illustrate the progressive changes that occur as neighborhood size increases. This specific choice (50, 150, 200) is justified by the observation that increasing the number of neighbors from 50 to 100 does not significantly increase the maximum distance between neighbors, nor does it result in noticeable changes in verifiability..
@@
We illustrate the progressive changes in fairness scores for the 8-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 17, 17 and 17.
@@
We illustrate the progressive changes in fairness scores for the 10-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 17, 17 and 17.
@@
Figure 12: Verifiability radar plot with gender-distance reference circle for 8-block with 50 nearest neighbors (experiment (4) in Figure 8, also in Section 4.2).
@@
Figure 15: Verifiability radar plot with gender-distance reference circle for 10-block with 50 nearest neighbors (experiment (7) in Figure 8, also in Section 4.2).
@@
C.3.4 Distances of words before and after pre-training
@@
It is important to note that not all related word pairs exhibited reduced distances after pre-training, nor did all unrelated pairs show increased separation. This is not a limitation of the pre-training process itself, but rather a consequence of computational constraints, which prevented exhaustive pairing of all possible related and unrelated words; doing so would have significantly expanded the data constructed for pre-training and potentially degraded the efficiency and performance of the embedding model [^14] Readers should note that the objective of this experiment (as well as for toxicity detection) is not to train the embedding model on an exhaustive set of related and unrelated word pairs, but rather to illustrate the effectiveness of the pre-training-based transfer learning approach in shaping semantically meaningful embeddings..
@@
Table 13: Top 50 most $\ell_{\infty}$similar words to female and male using the Keras embedding layer with pre-training.
Model Salary Prediction & Gender Bias Mitigation (8 & 10 blocks)
@@
Word Dist. ($\downarrow$) Word Dist. ($\downarrow$) Word Dist. ($\downarrow$) Word Dist. ($\downarrow$)
@@
Table 15: Top 50 most $\ell_{\infty}$similar words to bitch using the pre-trained Keras embedding layer.
Relation Similar words to bitch ($\ell_{\infty}$ distance)
[PAD] $1.645\text{e}{-03}$#tter $2.569\text{e}{-03}$ cock $2.631\text{e}{-03}$
@@
##kker $2.267\text{e}{-03}$#s $2.804\text{e}{-03}$#holz $2.854\text{e}{-03}$
##r $2.320\text{e}{-03}$ pack $2.856\text{e}{-03}$#1 $3.011\text{e}{-03}$
##z $2.409\text{e}{-03}$ wan $3.100\text{e}{-03}$#ie $3.137\text{e}{-03}$
bi $2.445\text{e}{-03}$#t $3.145\text{e}{-03}$ ; $3.154\text{e}{-03}$
##cks $2.533\text{e}{-03}$ $3.168\text{e}{-03}$#tt $3.179\text{e}{-03}$
##us $3.188\text{e}{-03}$#3 $3.261\text{e}{-03}$#ra $3.289\text{e}{-03}$
+ $3.340\text{e}{-03}$ fu $3.386\text{e}{-03}$#te $3.416\text{e}{-03}$
( $3.431\text{e}{-03}$#i $3.469\text{e}{-03}$#ot $3.471\text{e}{-03}$
##or $3.496\text{e}{-03}$#er $3.502\text{e}{-03}$ mother $3.514\text{e}{-03}$
@@
##dge $3.806\text{e}{-03}$#rac $3.837\text{e}{-03}$ ass $3.874\text{e}{-03}$
@@
sucker $3.976\text{e}{-03}$#ey $3.987\text{e}{-03}$#ore $4.006\text{e}{-03}$
ho $4.020\text{e}{-03}$ ku $4.116\text{e}{-03}$ $4.118\text{e}{-03}$
! $4.165\text{e}{-03}$#it $4.184\text{e}{-03}$
@@
A selection of randomly chosen synonym and antonym pairs, along with their semantic embedding distances before and after embedding pre-training, is presented in LABEL:tab:distance_metrics_gen_sen. As observed, synonym pairs exhibit reduced distances while antonym pairs become more distant in the embedding space after pre-training, demonstrating the intended effect of the contrastive learning approach.

### v2

@@ C.3 Ablation studies
Adversarial Attack
Robustness is a key prerequisite for reliable model verification, and adversarial training makes the models more robust (Goodfellow et al., 2015a; Bai et al., 2021); therefore, in this section, we examine whether our framework assigns higher fairness scores to models that undergo adversarial training. To better show the effect of adversarial training, we consider a model variant with embedding layers that differ from those used in Section 4.2. This design choice ensures that the conventionally trained model exhibits a fairness score of zero, providing a clear baseline model from which improvements introduced by adversarial training can be more clearly measured. Please note that the 2 models used for this ablation study have the exact same embedding layers so that we can isolate the effect of pretraining.
To improve robustness against perturbations, we apply offline adversarial data augmentation via synonym replacement. Given an existing labeled dataset, additional training samples are generated only for the positive class ($y=1$), while all original samples are retained. The number of variants per positive instance is controlled by an augmentation factor $k$. In our adversarial experiments, we set the augment factor to be 3, producing three additional augmented sentences for each original positive example. Detailed pretraining and training hyperparameters can be found in Table 11.
Each augmented sentence is created through token-level synonym substitution based on WordNet. For every non-numeric token in a sentence, synonym replacement is applied independently with probability $p$. We use a high replacement probability of $p=0.8$, which encourages the generation of more adversarial examples. This augmentation strategy introduces perturbations, thereby simulating adversarial paraphrases that increase the model’s robustness during training. The result can be found in Table 11.
Table 11: Training configuration, optimization strategy, and summary results for baseline and adversarially trained models. Fairness improvements are achieved through offline synonym-based augmentation during model training.
[…]
Model $\bm{\alpha}$ Margin $\bm{\alpha}$ (Gen. Sem.) Margin (Gen. Sem.) Epochs Epochs (Gen. Sem.) Aug. Prob. Aug. Factor
3 blocks (baseline) 20 0.3 1.0 0.3 100 100 0 0
3 blocks (adv. trained) 20 0.3 1.0 0.3 100 100 0.8 3
[…]
Model Vocab Seq. Len. Embed Dim Heads FF Dim Dropout Epochs Batch
3 blocks 30,522 200 8 2 8 0.2 150 128
[…]
Optimizer: AdamW (lr $10^{-3}$, weight decay $10^{-5}$)
LR scheduler: ReduceLROnPlateau (factor =0.5=0.5, patience =3=3, min lr $10^{-6}$)
Early stopping: Patience =10=10, restore best weights
Loss: Focal Loss ($\alpha=0.25$, $\gamma=2.0$)
[…]
Model Fairness Score Test Acc.
Baseline 0% 91%
Adversarially Trained 44% 91%
Minimum ($\inf$) -1.242537
Maximum ($\sup$) 1.161774
Range ($\sup-\inf$) 2.404311
Fairness scores without pre-training the Keras embedding layer
@@
The results for the 50 nearest neighbors and their corresponding distances for the words female and male are presented in Table 12, respectively. As shown, in the absence of pre-training, the nearest neighbors seldom reflect strongly gendered semantics. This underscores the importance of embedding layer pre-training in capturing meaningful semantic relationships, thereby justifying the design of our experimental framework. For comparison, the distance results with pre-training are provided in Section C.3, some examples of distances of strongly gender-related word pairs without specifically pre-training can be found in Table 13.
For the sake of completeness, we conducted the fairness score assessment despite the absence of gendered words among the 50 nearest neighbors for both gender indicators, we simplified the fairness evaluation by setting the threshold based solely on the $\ell_{\infty}$ distance between female and male (which is the gender indicator in the Adult Census dataset), as the reference circle (which is 0.7581), rather than using the maximum distance among the 50 neighboring words of the two gender indicators. Please note that the entire experimental setup, including the language model training parameters, remains identical to the experiment that includes the pre-training of the embedding model (Section 4.2 Experimental Results). The results of the formal verification for the model without any pre-training of the Keras embedding layer are presented in Figure 5. As shown, not only do all sentences fail verification, but their corresponding $\ell_{\infty}$ radii also lie significantly within the reference circle.
Table 12: Top 50 most $\ell_{\infty}$ -similar words to female and male using the Keras embedding layer without pre-training.
[…]
@@
Word Distance Word Distance Word Distance Word Distance
@@
7th $1.488\text{e}{-01}$ 6th $1.504\text{e}{-01}$ ##ec $1.465\text{e}{-01}$ specialty $1.474\text{e}{-01}$
5th $1.560\text{e}{-01}$ 8 $1.672\text{e}{-01}$ 85 $1.477\text{e}{-01}$ ##em $1.519\text{e}{-01}$
1st $1.753\text{e}{-01}$ 5 $1.794\text{e}{-01}$ 13 $1.534\text{e}{-01}$ ##v $1.568\text{e}{-01}$
@@
dominican $2.369\text{e}{-01}$ ##vi $2.384\text{e}{-01}$ england $2.080\text{e}{-01}$ lee $2.130\text{e}{-01}$
@@
Relationship between embedding space and max perturbed noise radius
This section investigates the relationship between the margin parameter in (2), the resulting embedding bounds, and the maximum verifiable perturbation radius. As the margin parameter decreases, the embedding bounds contract accordingly, leading to a reduction in the verifiable radius. This contraction can impede the successful verification of semantically similar words, as their embedding distances may fall outside the increasingly tightened bounds. An illustrative graph Figure 6 highlights the trade-off between margin size and verifiability within semantic clusters (The range in the graph is for demonstrative purposes only; for more detailed metrics corresponding to specific margin values, please refer to Section C.2).
@@
Transformer-neighbor configurations
In this section, we examine how fairness scores are influenced by varying combinations of transformer block counts and the number of nearest neighbors. The experimental configuration for this analysis is illustrated in Figure 7. As previously discussed in Section Experimental Results, we presented fairness scores for selected models with different transformer depths. Here, our objective is to provide a more comprehensive view, highlighting how fairness metrics evolve progressively as we adjust the number of transformer blocks and nearest neighbors. This analysis aims to uncover systematic trends and sensitivities within the model architecture that impact fairness.
@@
We illustrate the progressive changes in fairness scores for the 6-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 16, 16 and 16 (The selected numbers of neighbors are intended to illustrate the progressive changes that occur as neighborhood size increases. This specific choice (50, 150, 200) is justified by the observation that increasing the number of neighbors from 50 to 100 does not significantly increase the maximum distance between neighbors, nor does it result in noticeable changes in verifiability).
@@
We illustrate the progressive changes in fairness scores for the 8-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 16, 16 and 16.
@@
We illustrate the progressive changes in fairness scores for the 10-block salary prediction model, evaluated across different settings of nearest neighbors (50, 150, and 200). These results highlight the impact of neighborhood size on fairness evaluation and are visualized in Figures 16, 16 and 16.
@@
Figure 11: Verifiability radar plot with gender-distance reference circle for 8-block with 50 nearest neighbors (experiment (4) in Figure 7, also in Section Experimental Results).
@@
Figure 14: Verifiability radar plot with gender-distance reference circle for 10-block with 50 nearest neighbors (experiment (7) in Figure 7, also in Section Experimental Results.
@@
Distances of words before and after pre-training
@@
It is important to note that not all related word pairs exhibited reduced distances after pre-training, nor did all unrelated pairs show increased separation. This is not a limitation of the pre-training process itself, but rather a consequence of computational constraints, which prevented exhaustive pairing of all possible related and unrelated words; doing so would have significantly expanded the data constructed for pre-training and potentially degraded the efficiency and performance of the embedding model. Readers should note that the objective of this experiment (as well as for toxicity detection) is not to train the embedding model on an exhaustive set of related and unrelated word pairs, but rather to illustrate the effectiveness of the pre-training-based transfer learning approach in shaping semantically meaningful embeddings.
@@
Table 14: Top 50 most $\ell_{\infty}$ -similar words to female and male using the Keras embedding layer after pre-training.
[…]
@@
Word Distance ($\downarrow$) Word Distance ($\downarrow$) Word Distance ($\downarrow$) Word Distance ($\downarrow$)
@@
Table 16: Top 50 most $\ell_{\infty}$ -similar words to bitch using the pre-trained Keras embedding layer.
[…]
[PAD] $1.645\text{e}{-03}$ ##tter $2.569\text{e}{-03}$ cock $2.631\text{e}{-03}$
@@
##kker $2.267\text{e}{-03}$ ##s $2.804\text{e}{-03}$ ##holz $2.854\text{e}{-03}$
##r $2.320\text{e}{-03}$ pack $2.856\text{e}{-03}$ ##1 $3.011\text{e}{-03}$
##z $2.409\text{e}{-03}$ wan $3.100\text{e}{-03}$ ##ie $3.137\text{e}{-03}$
bi $2.445\text{e}{-03}$ ##t $3.145\text{e}{-03}$ ; $3.154\text{e}{-03}$
##cks $2.533\text{e}{-03}$ * $3.168\text{e}{-03}$ ##$t$ $3.179\text{e}{-03}$
[…]
+ $3.340\text{e}{-03}$ fu $3.386\text{e}{-03}$ ##te $3.416\text{e}{-03}$
( $3.431\text{e}{-03}$ ##i $3.469\text{e}{-03}$ ##ot $3.471\text{e}{-03}$
##or $3.496\text{e}{-03}$ ##er $3.502\text{e}{-03}$ mother $3.514\text{e}{-03}$
@@
##dge $3.806\text{e}{-03}$ ##rac $3.837\text{e}{-03}$ ass $3.874\text{e}{-03}$
@@
sucker $3.976\text{e}{-03}$ ##ey $3.987\text{e}{-03}$ ##ore $4.006\text{e}{-03}$
ho $4.020\text{e}{-03}$ ku $4.116\text{e}{-03}$ - $4.118\text{e}{-03}$
! $4.165\text{e}{-03}$ ##it $4.184\text{e}{-03}$
@@
A selection of randomly chosen synonym and antonym pairs, along with their semantic embedding distances before and after embedding pre-training, is presented in Table 17. As observed, synonym pairs exhibit reduced distances while antonym pairs become more distant in the embedding space after pre-training, demonstrating the intended effect of the contrastive learning approach.